OpenStudy (anonymous):

I WILL GIVE A MEDAL AND FAN YOU! Factor Completely: 12a3d2 − 6ad3. Prime 6a3d3(2a − d) 6ad2(2ad − d) 6ad2(2a2 − d)

OpenStudy (anonymous):

Is this your question? \(12a^3d^2 - 6ad^3\)

OpenStudy (anonymous):

Yes Ma'am

OpenStudy (anonymous):

So, look for common terms. Start with the coefficients.

OpenStudy (anonymous):

I meant common factors, sorry.

OpenStudy (anonymous):

So like the 6a and the 12a?

OpenStudy (anonymous):

Yes, but you will be dividing so only use the terms that will divide evenly.

OpenStudy (anonymous):

First look at 6.

OpenStudy (anonymous):

Can you divide both terms by 6?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Its the last one:

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

I know that i can divide it by 12 by 6

OpenStudy (anonymous):

yw!

OpenStudy (anonymous):

So divide both terms by 6. Look at what you have left and see what else you can divide both terms by.

OpenStudy (anonymous):

which ones are the terms :(

OpenStudy (anonymous):

One term is \(12a^3d^2\) and the other is \(6ad^3\)

OpenStudy (anonymous):

Ohh So dividing 12/6 and divide 6/6

OpenStudy (anonymous):

Yes. Then look at the variables and see what else you can factor out.

OpenStudy (anonymous):

well there are two of these ^3

OpenStudy (anonymous):

\(6(2a^3d^2 - ad^3)\)

OpenStudy (anonymous):

It is the variables you are looking at, not their exponents. You cannot factor out an exponent separate from its variable.

OpenStudy (anonymous):

ohh so there are two a's and two d's

OpenStudy (anonymous):

sorry i thought i been entered that

OpenStudy (anonymous):

Each term can be divided by a and by d, the question is how many of each can be factored out.

OpenStudy (anonymous):

Sorry, you did mention the a earlier. I must have missed it.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

when they say factor , what does that stand for?

OpenStudy (anonymous):

2 and 3 are factors of 6. \(2 \times 3 = 6\) Factors are the numbers (or variables) that multiply together to make a product(the answer of multiplication).

OpenStudy (anonymous):

Ohhh

OpenStudy (anonymous):

So, You've factored out a 6 and now you want to factor out an a. What does that leave you with inside the parenthesis?

OpenStudy (anonymous):

2a^2-d?

OpenStudy (anonymous):

\(6(2a^3d^2 - ad^3)\) Only 1 a can be factored out so divide each term by a. Remember when you are dividing the same base with different exponents, you will subtract the exponents.

OpenStudy (anonymous):

Ohh well that makes more sense

OpenStudy (anonymous):

So, \(\Large{\frac{2a^3d^2}{a}} = \)what?

OpenStudy (anonymous):

Then do the same with the other term.

OpenStudy (anonymous):

how could they be divided by a variable?

OpenStudy (anonymous):

All the numbers and variables in the term are being multiplied so when you divide them, it is just like canceling numbers. Let me try to demonstrate: \[\frac{4\times2 \times 3}{3 \times 6 \times 2}\] You can divide the 2 by 2 and the 3 by 3 and end up with \(\frac{4}{6}\).

OpenStudy (anonymous):

Ohh and just be left with 4/6

OpenStudy (anonymous):

but what about the bottom?

OpenStudy (anonymous):

its only a variable

OpenStudy (anonymous):



OpenStudy (anonymous):

Do you see what is left?

OpenStudy (anonymous):

yes 2a^2d^2

OpenStudy (anonymous):

like that?

OpenStudy (anonymous):

Ok, now do the same to the other term.